OpenStudy (anonymous):

Can someone please help me simplify this.. 6x^2 v^9*2y^-1 x^-9*5v^-9 y

OpenStudy (compassionate):

Hello, do you know the first step to simplifying?

OpenStudy (whpalmer4):

\[6x^2v^9*2y^{-1}x^{-9}*5v^{-9}y\] Correct?

OpenStudy (whpalmer4):

I would recombine that as \[(6*2*5)*(v^9*v^{-9})*(x^2*x^{-9})*(y^{-1}*y)\] Use the law of multiplication for exponents with common bases: \[u^a*u^b = u^{a+b}\]to do some more simplification. Don't forget that \(u^1 = u\) and \(u^0 = 1\)

OpenStudy (anonymous):

Would my answer be \[\frac{ 60 }{ x ^{7}y} \]

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

Not quite...

OpenStudy (whpalmer4):

6*2*5 = 60, that part is right the v terms combine to = 1 the x terms combine to = x^-7 have another look at the "Don't forget" comment when contemplating the value of the y terms